An alpha particle (2 protons + 2 neutrons) is brought to a stop by a potential difference of 450 V . What was its initial speed?


Express your answer using two significant figures.
1
Expert's answer
2019-09-27T09:33:07-0400

The work done to stop the particle is "A = q V". Since the mechanical energy is conserved, "T_i - A = T_f", where "T_{i,f}" is the kinetic energy of alpha particle in the initial and final state: "T_i = \frac{m v_i^2}{2}" , "T_f = 0".

The charge of the alpha particle is "q = 2 e = 3.2 \cdot 10^{-19} C" , and the mass of the alpha particle is "6.64 \cdot 10^{-27} kg".

Hence, the initial speed was "v_i = \sqrt{\frac{2 A}{m}} = \sqrt{\frac{2 q V}{m}} = 208.26 \cdot 10^6 \frac{m}{s}".